# Benchmarking Expressive Japanese Character Text-to-Speech with VITS and Style-BERT-VITS2

### v1

@@ 5.1 Mean Opinion Score
All models performed comparably to the ground truth in terms of naturalness. Hikari TTS achieved a score of $4.47\pm 0.68$, slightly underperforming relative to its ground truth counterpart ($4.76\pm 0.48$). This indicated that raters generally perceived the ground truth audio as more natural than the TTS-generated audio for Hikari. Conversely, Kitsune TTS slightly outperformed its ground truth, with scores of $4.40\pm 0.78$ and $4.14\pm 0.89$, respectively. This suggested that raters found the TTS-generated audio to be more natural than the ground truth for Kitsune, indicating that certain aspects of the synthesized audio surpassed the quality of the ground truth recordings. Finally, Saya TTS and Saya ground truth were nearly identical, receiving scores of $4.24\pm 0.76$ and $4.24\pm 0.75$, respectively, suggesting parity in perceived naturalness. Kitsune ground truth exhibited the highest variability (±$\pm 0.89$), despite being recorded by a professional voice actor. This highlighted subjective differences in rater perceptions. Moreover, the overall score for TTS ($4.37\pm 0.74$) was nearly identical to the overall ground truth score ($4.38\pm 0.77$). A two-tailed, paired t-test on aggregate rating data yielded an insignificant result with $p\approx 0.91$, indicating that native speakers did not perceive a statistically significant difference in naturalness between the SBV2JE audio and the ground truth recordings.
5.2 Comparative Mean Opinion Score
To evaluate the difference in human preference between VITS and SBV2JE, we used Comparative Mean Opinion Score with 5 native Japanese speakers as raters. Raters were presented two audios generated using the same text, one from VITS and one from SBV2JE, and were asked to rate their preference on a scale of -3 (greatly prefer the second model) to 3 (greatly prefer the first model). The position of models presented was randomized in addition to the speakers. 10 audios were used from each speaker, resulting in a total of 150 ratings.
Table 2: CMOS Results for TTS Models
@@ 5.2 Comparative Mean Opinion Score
5.3 Intra-Speaker Consistency
We calculated intra-speaker similarity by leveraging speaker embeddings generated by the pretrained Resemblyzer model. Audio files were grouped by speaker and model, and each file was processed to extract embeddings representing the speaker’s identity. For each speaker-model pair, we computed cosine similarity between all pairs of embeddings, which quantified how consistently the model maintained the same speaker’s identity across multiple audio samples. The average similarity score for each speaker-model pair was calculated, and these scores were further grouped by speaker and model to evaluate trends. This approach provided a robust measure of how well each TTS model preserved speaker consistency.
Table 3: Intra-Speaker Similarity Scores for TTS Models
@@ 5.3 Intra-Speaker Consistency
Both the VITS and SBV2JE models achieved high intra-speaker consistency across all speakers, with similarity scores exceeding 95% in all cases. Among the speakers, Kitsune exhibited the highest average similarity score (97.05%), indicating that the models handled this speaker’s identity particularly well. In contrast, Saya showed slightly lower consistency (95.78%), suggesting that maintaining this speaker’s identity was more challenging for the models. When grouped by model, VITS slightly outperformed SBV2JE (96.43% vs. 96.24%), although the difference was minimal. These results highlighted the robustness of both models in preserving speaker identity, with only minor variations across speakers and models. This strong performance was encouraging for real-world applications where consistent speaker representation was essential, such as in character-driven TTS systems for language learning. Further analysis could explore whether differences in the training data or speaker characteristics contributed to the observed variability.
5.4 Word Error Rate
We calculated Word Error Rate to assess intelligibility by comparing generated speech against corresponding ground truth text. First, we transcribed the audio files using the Faster-Whisper ASR model, initialized with a seed prompt for better context. To ensure linguistic accuracy, we tokenized both the transcriptions and ground truth text using Sudachi, a Japanese morphological analyzer Takaoka et al. (2018). For each transcription, we identified the most relevant ground truth line using Levenshtein similarity matching, ensuring a fair comparison even if minor differences existed between the audio content and ground truth. Finally, WER was calculated by measuring the proportion of word-level errors (insertions, deletions, and substitutions) between the tokenized transcription and the matched ground truth. This methodology provided a robust evaluation of how intelligible and accurate the TTS outputs were when processed through a real-world ASR system.
Table 4: Word Error Rate Comparison
@@ 5.4 Word Error Rate
Both the VITS and SBV2JE models achieved high intelligibility across all speakers, with consistently low WER values. "Kitsune" showed the best performance, with SBV2JE achieving the lowest WER (0.02), while "Saya" had the highest WER (0.08) for both models. When grouped by model, SBV2JE slightly outperformed VITS in overall WER (0.04 vs. 0.05) with comparable variability. These results highlighted both models’ reliability, with SBV2JE offering slightly greater consistency, making them well-suited for character-driven TTS applications in language learning.
6 Discussion
@@
While these improvements make SBV2JE well-suited for character-driven conversational agents and language learning applications, the model can be applied to storytelling, anime voiceover, personalized voice assistants, and other applications where naturalness and expressiveness are critical. While untested in this study, the model’s control over stylistic elements could open up opportunities for more adaptive and engaging user experiences in interactive platforms. However, there are limitations to consider. For example, SBV2JE presents trade-offs between model complexity and computational efficiency. While the model’s enhancements yield measurable gains, the increased computational cost associated with its architecture, such as higher gin channel values and WavLM-based discriminators, could pose challenges for deployment on resource-constrained devices.
@@
7 Conclusion
SBV2JE represents a significant advancement in Japanese TTS character language synthesis, balancing naturalness, intelligibility, and expressiveness. By addressing linguistic nuances and leveraging innovative architectural modifications, it sets a benchmark for character-driven speech synthesis and language learning. Future research should aim to refine its generalization, explore new use cases, and optimize its deployment to maximize its impact across diverse applications.

### v2

@@ V-A Mean Opinion Score
All models performed comparably to the ground truth in terms of naturalness. Hikari TTS achieved a score of $4.47\pm 0.68$, slightly underperforming relative to its ground truth counterpart ($4.76\pm 0.48$). This indicated that raters generally perceived the ground truth audio as more natural than the TTS-generated audio for Hikari. Conversely, Kitsune TTS slightly outperformed its ground truth, with scores of $4.40\pm 0.78$ and $4.14\pm 0.89$, respectively. This suggested that raters found the TTS-generated audio to be more natural than the ground truth for Kitsune, indicating that certain aspects of the synthesized audio surpassed the quality of the ground truth recordings. Finally, Saya TTS and Saya ground truth were nearly identical, receiving scores of $4.24\pm 0.76$ and $4.24\pm 0.75$, respectively, suggesting parity in perceived naturalness. Kitsune ground truth exhibited the highest variability ($\pm 0.89$), despite being recorded by a professional voice actor. This highlighted subjective differences in rater perceptions. Moreover, the overall score for TTS ($4.37\pm 0.74$) was nearly identical to the overall ground truth score ($4.38\pm 0.77$). A two-tailed, paired t-test on aggregate rating data yielded an insignificant result with $p\approx 0.91$, indicating that native speakers did not perceive a statistically significant difference in naturalness between the SBV2JE audio and the ground truth recordings.
V-B Comparative Mean Opinion Score
To evaluate the difference in human preference between VITS and SBV2JE, we used Comparative Mean Opinion Score with 5 native Japanese speakers as raters. Raters were presented two audios generated using the same text, one from VITS and one from SBV2JE, and were asked to rate their preference on a scale of -3 (greatly prefer the second model) to 3 (greatly prefer the first model). The position of models presented was randomized in addition to the speakers. Ten audios were used from each speaker, resulting in a total of 150 ratings.
TABLE II: CMOS Results for TTS Models
@@ V-B Comparative Mean Opinion Score
V-C Intra-Speaker Consistency
We calculated intra-speaker similarity by leveraging speaker embeddings generated by the pretrained Resemblyzer model [Resemblyzer]. Audio files were grouped by speaker and model, and each file was processed to extract embeddings representing the speaker’s identity. For each speaker-model pair, we computed cosine similarity between all pairs of embeddings, which quantified how consistently the model maintained the same speaker’s identity across multiple audio samples. The average similarity score for each speaker-model pair was calculated, and these scores were further grouped by speaker and model to evaluate trends. This approach provided a robust measure of how well each TTS model preserved speaker consistency.
TABLE III: Intra-Speaker Similarity Scores for TTS Models
@@ V-C Intra-Speaker Consistency
Both the VITS and SBV2JE models achieved high intra-speaker consistency across all speakers, with similarity scores exceeding 95% in all cases. Among the speakers, Kitsune exhibited the highest average similarity score (97.05%), indicating that the models handled this speaker’s identity particularly well. In contrast, Saya showed slightly lower consistency (95.78%), suggesting that maintaining this speaker’s identity was more challenging for the models. When grouped by model, VITS slightly outperformed SBV2JE (96.43% vs. 96.24%), although the difference was minimal. These results highlighted the robustness of both models in preserving speaker identity, with only minor variations across speakers and models. This strong performance is encouraging for real-world applications where consistent speaker representation is essential, such as in character-driven TTS systems for language learning. Further analysis could explore how differences in speaker characteristics can contribute to variability.
V-D Word Error Rate
We calculated Word Error Rate to assess intelligibility by comparing generated speech against corresponding ground truth text. First, we transcribed the audio files using the Faster-Whisper ASR model [fasterwhisper], initialized with a seed prompt for better context. To ensure linguistic accuracy, we tokenized both the transcriptions and ground truth text using Sudachi, a Japanese morphological analyzer [TAKAOKA18.8884]. For each transcription, we identified the most relevant ground truth line using Levenshtein similarity matching, ensuring a fair comparison even if minor differences existed between the audio content and ground truth. Finally, WER was calculated by measuring the proportion of word-level errors (insertions, deletions, and substitutions) between the tokenized transcription and the matched ground truth. This methodology provided a robust evaluation of how intelligible and accurate the TTS outputs were when processed through a real-world ASR system.
TABLE IV: Word Error Rate Comparison
@@ V-D Word Error Rate
Both the VITS and SBV2JE models achieved high intelligibility across all speakers, with consistently low WER values. Kitsune showed the best performance, with SBV2JE achieving the lowest WER (0.02), while Saya had the highest WER (0.08) for both models. When grouped by model, SBV2JE slightly outperformed VITS in overall WER (0.04 vs. 0.05) with comparable variability. These results highlighted both models’ reliability, with SBV2JE offering slightly greater consistency, making them well-suited for character-driven TTS applications in language learning.
VI Discussion
@@
While these improvements make SBV2JE well-suited for character-driven conversational agents and language learning applications, the model can be applied to storytelling, anime voiceover, personalized voice assistants, and other applications where naturalness and expressiveness are critical.
In addition to their use in anime-focused language learning tools, expressive TTS models like VITS and Style-BERT-VITS2 are well-suited for live storytelling, where emotional nuance plays a key role in delivering a compelling narrative. Although our work focuses on Japanese, both models support English and Chinese as well, opening the door to cross-lingual applications such as voice translation and code-switching across these languages. These models can also be deployed as backend APIs for mobile or edge devices, broadening their practical utility across platforms. Finally, speech synthesis connects users through the Internet of Things (IoT) [zhao2023emotion] and personalized voice assistance.
However, there are limitations to consider. For example, SBV2JE presents trade-offs between model complexity and computational efficiency. While the model’s enhancements yield measurable gains, the increased computational cost associated with its architecture, such as higher gin channel values and WavLM-based discriminators, could pose challenges for deployment on resource-constrained devices.
While untested in this study, the model’s control over stylistic elements could open up opportunities for more adaptive and engaging user experiences in interactive platforms. These stylistic elements open future work in fine-grained emotion control, real-time conversational adaptation, cross-lingual expressive synthesis, personalized voice tutoring for language learners, and low-resource style transfer for underrepresented character voices.
Our study is constrained by the small per-character datasets, the limited number of characters and raters, and the SBV2JE repository’s use of all available audio for training, and while these factors restrict generalization and efficiency analysis, our evaluations focus on perceptual judgments of generated speech, leaving broader validation to future work.
@@
VII Conclusion
SBV2JE represents a significnt advancement in Japanese TTS character language synthesis, balancing naturalness, intelligibility, and expressiveness. By addressing linguistic nuances and leveraging innovative architectural modifications, it sets a reference standard for character-driven speech synthesis and language learning. Future research should aim to refine its generalization, explore new use cases, and optimize its deployment to maximize its impact across diverse applications.